# Data-driven methods to discover stable linear models of the helicity injectors on HIT-SIU

### v1

@@ Abstract
Accurate and efficient circuit models are necessary to control the power electronic circuits found on plasma physics experiments. Tuning and controlling the behavior of these circuits is inextricably linked to plasma performance. Linear models are greatly preferred for control applications due to their well-established performance guarantees, but they typically fail to capture nonlinear dynamics and changes in experimental parameters. Data-driven system identification can help mitigate these shortcomings by learning interpretable and accurate reduced-order models of a complex system, in this case the injector circuits of the Helicity Injected Torus – Steady Inductive Upgrade (HIT-SIU) experiment. Specifically, the Bagging Optimized Dynamic Mode Decomposition (BOP-DMD), is leveraged to learn stable, reduced order models of the interaction between the spheromak plasma formed in the confinement volume, and the injector circuits of the device. BOP-DMD is trained and evaluated on an analytic model of the vacuum dynamics of the injector circuits of HIT-SIU, as well as an analytic linear reduced-order model for the injector dynamics when a plasma is present. BOP-DMD is then fit on experimental data, both on shots with and without a plasma in the confinement volume. In doing so, we demonstrate the capability of data-driven methods to produce stable, linear models for control and uncertainty quantification in plasma experiments.
@@ I Introduction:
While great progress has been made in the last few decades towards viable nuclear fusion reactors [33, 12], the problem of controlling the multi-scale and dynamic plasma formed in the confinement volumes of these devices remains an active area of research [16, 5, 44, 39]. Many modern control algorithms require a model of the dynamics of interest [3], or rely on a black-box, model-free control strategy [11]. Moreover, while there are a variety of models for describing the evolution of a plasma under fusion conditions, many of the parameters upon which these equations depend are difficult to measure in modern experiments, or exhibit highly nonlinear behavior, making them difficult to incorporate into a real-time control scheme. However, the data-driven method of dynamic mode decomposition advocated here, which is a regression to a best-fit linear model, is demonstrated to provide a viable path towards stable control of plasmas.
@@
The diagnostics, circuits, and power supplies utilized on modern plasma physics experiments provide a wealth of data to which modern system identification techniques can be applied to learn the underlying dynamics of a particular interaction [23]. In particular, linear reduced order models can be learned from time-series data, and then implemented in optimal linear feedback control and estimation. Techniques such as the Dynamic Mode Decomposition (DMD) [42] allow for the discovery of such models, while also providing an interpretable (i.e. with explicit equations) representation of the dynamics. Optimal control and estimation has also been popular in plasma physics for some time, with examples in tokamak control [30, 6] being a particular successful application of these techniques.
@@
In recent years, there has been work showing the promise of pairing interpretable data-driven system identification techniques with control. Nonlinear system-identified plasma models have been built for a number of configurations, the L-H transition in tokamaks [8], MHD simulations [27], kinetic simulations [2], plasma closures [48], turbulence modelling [1], 2D electrostatic drift-wave turbulence [15], divertor detachment in 2D simulations [32], and quasi-axisymmetric stellarators [43]. In addition, linear data-driven models have been built using the DMD for predicting the plasma dynamics in the HIT-SI line of devices [47, 26], kinetics prediction and acceleration [37, 36, 38], and learning the dynamics of $E\times B$ plasmas [13, 14]. Nonlinear and high-dimensional data-driven models with control, often using deep reinforcement learning, have also been successfully applied in recent years to learn control policies for tokamak shape design [11] and instability control [45].
@@
In this work we focus on modeling and controlling the injector circuits of the Helicity Injected Torus – Steady Inductive Upgrade (HIT-SIU), which are responsible for forming and sustaining the spheromak plasma inside the vacuum chamber. While these circuits exhibit a nonlinear coupling to the plasma, past work has shown that this interaction can be well approximated by a linear state space model  [29]. Further, we are interested in being able to easily modify an identified dynamics model without an expensive retraining phase, whether that be in terms of computational cost or accumulated training data. As such, the goal of this work is to implement the bagging optimized dynamic mode decomposition (BOP-DMD) [4, 41] to discover linear models for the interaction between spheromak plasmas formed in HIT-SIU, and the helicity injectors used to form and sustain the plasma. BOP-DMD has a number of advantages over standard DMD, including (i) being optimally robust to noise, (ii) providing uncertainty quantification metrics, and (iii) allowing for constraining the linear model to be stable. BOP-DMD models are then paired with optimal control and estimation, in the form of linear quadratic Gaussian control (LQG), with the objective of controlling the current profiles of the flux and voltage coils on the helicity injectors.
@@
I.1 HIT-SIU
The HIT-SIU experiment is studied using steady inductive helicity injection (SIHI) [25, 49] to form and sustain spheromak plasmas. Energy and helicity are injected into the plasma using a set of four injectors, each comprised of a voltage and a flux coil that link a semi-toroidal channel that connects to the main plasma volume. In each injector the flux coil injects, locally toroidal, magnetic flux, while the voltage coil acts as an air core transformer generating a voltage parallel to the flux-coil-generated field. The magnetic fields of the voltage and flux coils are orthogonal to one another such that there is no mutual inductance between the two coils when no plasma is present. However, all four injectors are part of common manifold structure (Fig. LABEL:fig:hitsiu-CAD), so the four voltage coils, and separately all four flux coils, are inductively coupled. Each coil is part of a larger resonant circuit comprised of an additional inductor and capacitor, which is driven by a H-bridge switching power amplifier (SPA). The SPAs drive a three-state square wave with a high, low, and zero crossing time. An example power supply waveform is shown in Fig. 2. The injector circuit topology is shown on the left side of Fig. 4. While the voltage and flux circuits have the same topology, the values of each individual element differ between the two circuits. However, both circuits are tuned to the same resonant frequency.
@@ I.1 HIT-SIU
I.2 Injector Circuit Control Methods:
Fast, and accurate control methods are paramount for creating, and sustaining, high performance plasmas in present-day and future fusion devices. The shape of the voltage and current waveforms for the HIT-SIU voltage and flux coils can determine the kinds of instabilities induced in the spheromak, as well as the overall helicity injected into the spheromak. All coils of the HIT-SIU experiment were controlled using a GPU-based proportional integral derivative (PID) controller [35]. The PID algorithm parameterized the target waveforms to be driven as a sinusoid with an amplitude $V_{0}$ or $\Psi_{0}$, frequency $f_{\text{inj}}$, phase $\phi_{A},\phi_{B},\phi_{C},\phi_{D}$, and offset, and used these control parameters to drive the desired current waveform through the coil. If the frequency and amplitude of the waveform driven on each injector is the same, the expression for the rate of helicity injection is given by
@@ I.2 Injector Circuit Control Methods:
While the PID algorithm implemented in  [35] was a significant improvement upon the previous control scheme for the HIT-SI devices, further improvements can be made by shifting from model-free control to model-based control. Linear optimal control is the most common kind of model-based control, and has been utilized across science and engineering  [21].
@@
Optimal linear control involves pairing an optimal state estimator, which reconstructs the full state of the system from limited noisy measurements, with a feedback controller, which calculates the optimal input to the system to track a desired trajectory. This optimal estimator, when constructed from a solution to the algebraic Riccatti equation, is known as a Kalman filter, and the feedback controller, when also found as the solution to the Riccatti equation, is called the linear quadratic regulator (LQR). These methods will pair together for LQG control. Both the Kalman filter and LQR rely on an underlying linear dynamics model of the system so that the current state and optimal gain to apply can be computed.
Figure 3: A representative equilibrium, with current gain of 6, showing field lines linking (gray) and isolated (rainbow) from the injector volume.
@@
Figure 4: Circuit topology for a voltage or a flux circuit that for a single injector. This diagram includes the SPA (S1-S4), the series inductor (L1), a parallel capacitance (C), and a parallel inductance (L2) which represents the voltage or the flux coil. a) is an example of a flux or voltage circuit in vacuum, and b) is flux or voltage circuit when a plasma is present. Lp denotes the inductance of the plasma, Rp it’s resistance, and Mp the mutual inductance between the circuit and the plasma.
II Analytic Model
@@
While there have been studies [28, 18, 19, 22] on the injector dynamics of the HIT-SI family of devices using MHD simulations and nonlinear reduced order models (ROMs), there has been less work on representing the interaction between the plasma and the injector circuits for the purposes of control. Before this study, the work of  [29], which derived a linear state space model of the injector circuit dynamics on HIT-SI3 and implemented a Kalman filter, was the only significant effort. To derive the linear model that will form the basis for the LQG control loop we seek to implement, we will begin with the circuit topology for the injector circuits as shown in Fig.  4. Treating the voltage across the capacitor, and the current through each coil as the states of our system, and then incorporating the mutual inductance between each flux or voltage coil, yields a 12-dimensional coupled system. This 12-dimensional system will be referred to as the vacuum circuit model. When a plasma is present one or more additional states need to be added to the model. In this work we choose to represent the plasma as a series LR circuit, which is coupled with a mutual inductance to each voltage or flux circuit. This yields a 13-dimensional coupled linear model. The full analytic vacuum model is available in Appendix A.
@@
The values used in the state space model for the inductance’s of the various circuit elements were obtained by applying a Hilbert transform to the voltage and current waveform from each element during a vacuum shot. However, there is a slight difference between the capacitance used in the circuit model, and the listed capacitance for the flux circuits in HIT-SIU. The capacitance’s used in the model range from 94.08 to 123.96 microfarads. These values were chosen by inspection to achieve better matching between the phases of the flux coil current waveforms of the model, and the observed flux coil current waveforms during vacuum shots.
@@
This model does make some important assumptions. Specifically, we assume that when a plasma is present in the confinement volume, there is no inductive coupling between the plasma and the series inductor in each injector circuit. This assumption, which is well motivated as the series inductor is separated by a significant distance from the device, then allows us to conclude that the matrix that determines how actuation affects the dynamics of the injector circuits is constant between vacuum and plasma regimes. This can be observed from the structure of the $\mathbf{B}$ matrix shown in Appendix VI. By using standard state space techniques for circuit analysis, it’s clear that the power supply input will only directly change the state of the series inductor. Thus the structure of the circuit is such that we can isolate differences between the vacuum and plasma regimes to the flux or voltage coils.
@@
Figure 6: The analytic linear model for the flux circuits when a plasma is present is simulated using a power supply waveform from the ninth shot on August 16th, 2022. The analytic model represents the spheromak as an inductor that is mutually coupled to the other flux circuits. To the right of the vertical blue line, there is no plasma present, and to the right, a plasma is present.
II.1 Matlab and Simulink Model:
Developing control methods for hardware first requires a demonstration of performance in a digital environment. To that end, a model of the coupled flux circuits with full feedback control was developed in Matlab and Simulink. To verify the impedance’s of the various circuit elements, shot data was analyzed to capture the reactance and resistance of the elements, and to capture the strong and weak mutual inductance’s present between the injector circuits. Since the vacuum model can be understood as a special case of the plasma model with no mutual inductance between the plasma and the flux coils, we use the same impedance for the flux circuit elements across models. However, the value chosen for the impedance of the plasma in the plasma model, the resistance in the plasma, and the coupling between the plasma and the flux coils, were chosen by inspection, and intuition about the behavior of the plasma. To evaluate the performance of the analytic vacuum and plasma models with respect to capturing the dynamics present during an experimental shot, an SPA waveform from a shot is used as inputs to the model, and the outputs of the model are evaluated against the flux circuit measurements from the experiment. The results of this model are shown in Fig. 5. and Fig. 6 for the vacuum and plasma models respectively. While the analytic model is able to capture much of the behavior of the flux circuits in vacuum, the plasma model struggles to capture the behavior of the flux circuits when a plasma is present. Fig. 7.
@@ II.1 Matlab and Simulink Model:
With the derivation of a model that accurately captures the dynamics of interest in our system, the next step in the control scheme is to implement a Kalman filter. The Kalman filter is able to accurately reconstruct the true state of the system at a given time from noisy, limited measurements. Fig. 8 demonstrates the performance of the Kalman filter when reconstructing the voltage across the capacitor in the first flux circuit. While the filter only has access to four noisy measurements of the flux coil currents, it is still able to accurately predict the value of the capacitor voltage throughout the shot.
@@
Finally, we contend with the controller: LQR minimizes a cost function $J$, that weighs the cost of control, with the value of producing an accurate reconstruction of the desired trajectory. These two costs are weighted by hyperparameters that can be tuned to prioritize accuracy of a specific state, and in this case was picked to be the current through the flux or voltage coil, as this trajectory is most closely linked to plasma performance. In this paradigm, a Kalman filter is used to predict from limited noisy measurements the value of each state of the system at a given time. The LQR controller then makes a control decision to push the system to the desired state. The full control loop for this scenario is shown in Fig. 9.
@@
First, one must decide what the desired trajectory for the system is. HIT-SIU has three primary operating regimes that differ in the relative temporal phasing of the current in each flux coil. For this simulation, we choose the case where all circuits have the same temporal phasing, which is the simplest case of operation on HIT-SIU. The desired current waveform is arrived at by simulating the analytic model of the circuits with a standard power supply waveform seen on experimental shots. This desired waveform is then subtracted off from the flux circuit predictions that are the outputs of the Kalman filter. The LQR controller gain is then applied to the full state estimation with the desired trajectory subtracted off, allowing for trajectory tracking, instead of fixed point tracking. A power supply waveform generated by the LQR controller is shown in Fig. 10. The action of the controller on the system produces a current profile through the flux circuit that closely matches the desired profile Fig. 11.
@@
Figure 8: The Kalman filter has access to four noisy flux coil currents, and none of the other eight states, and is able to reconstruct the remaining states of the system. Shown here is the reconstruction of the capacitor voltage from the first flux circuit.
@@
Figure 9: Full LQG feedback control loop in Simulink. The Kalman filter receives as input the states corresponding to the current through the flux coils, and from this measurement reconstructs the full state of the system. The desired flux coil current profiles are then subtracted off from the current state of flux coils, and fed into the LQR controller.
@@
Figure 10: Voltage waveform determined by the LQR controller to best track the desired flux coil waveform.
@@
Figure 11: Output of the flux coil very closely matches the desired current waveform.
III Role of Data Driven Models

### v3

# Data-driven methods to discover stable linear models of the inductive helicity injectors

@@ Abstract
Accurate and efficient circuit models are necessary to control the power electronic circuits employed in most plasma physics experiments. Controlling the behavior of these circuits is inextricably linked to generating the desired plasma conditions. Linear models are greatly preferred for control applications due to their well-established performance guarantees, however, it is challenging to identify a linear approximation of the coupled nonlinear coil-plasma system from limited noisy measurements. In this work, the Bagging Optimized Dynamic Mode Decomposition (BOP-DMD) is leveraged to learn stable, reduced order models of the interaction between the coils and the plasma in the Helicity Injected Torus – Steady Inductive Upgrade (HIT-SIU) experiment. BOP-DMD is trained and evaluated on an analytic model of the vacuum dynamics of the injector circuits, as well as an analytic linear reduced-order model for the circuit dynamics when a plasma is present. BOP-DMD is then fit on experimental data, both on experiments with and without a plasma. In doing so, we demonstrate the capability of BOP-DMD to produce stable, linear models for control and uncertainty quantification in a high-power, coupled plasma-coil system.
@@ I Introduction:
Forming and sustaining a plasma for experimental or applied purposes generally requires the use of applied electric/magnetic fields to ionize/heat the plasma, drive currents, and/or maintain or modify embedded magnetic fields. This creates a coupled plasma-circuit system that can introduce significant nonlinearity and additional couplings between circuit elements that are not present in vacuum. Accurate and robust control of such systems is required to achieve the desired plasma conditions, and to ensure that the device remains in safe operational limits. The problem of controlling the multi-scale and dynamic plasma formed in experimental and industrial devices remains an active area of research [17, 1, 49, 43, 6, 15]. Moreover, the models and algorithms employed in real-time control should be interpretable, and have quantifiable uncertainty. Many modern control algorithms require a model of the dynamics of interest [4], or rely on a black-box, model-free control strategy [12]. Moreover, while there are a variety of models for describing the evolution of a plasma, many of the parameters upon which these equations depend are difficult to measure in modern experiments, or exhibit highly nonlinear behavior, making them difficult to incorporate into a real-time control scheme. However, the data-driven method of dynamic mode decomposition advocated here, which is a regression to a best-fit linear model, is demonstrated to provide a viable path towards stable control of the coupled coil plasma system.
@@
The diagnostics, circuits, and power supplies utilized on modern plasma physics experiments provide a wealth of data to which modern system identification techniques can be applied to learn the underlying dynamics of a particular interaction [25]. In particular, linear reduced order models can be learned from time-series data, and then implemented in optimal linear feedback control and estimation. Techniques such as the Dynamic Mode Decomposition (DMD) [48] allow for the discovery of such models, while also providing an interpretable (i.e. with explicit equations) representation of the dynamics. Optimal control and estimation has also been popular in plasma physics for some time, with examples in tokamak control [34, 7] being a particularly successful application of these techniques, as well as in plasma processing for semiconductor manufacturing [46, 45].
@@
In recent years, there has been work showing the promise of interpretable data-driven system identification techniques in plasma physics. Nonlinear system-identified plasma models have been built for a number of complex plasma systems/dynamics, including: MHD simulations of helicity injection [29], kinetic plasma simulations [3], plasma closures [52], plasma sources for semiconductor etching [32], 2D electrostatic drift-wave turbulence [16], turbulence bifurcations in fusion systems [9], coupled plasma-neutral systems [37], and ablative pulsed plasma thrusters [24]. In addition, linear data-driven models have been built using the DMD for predicting the plasma dynamics in the HIT-SI line of devices [51, 28], kinetics prediction and acceleration [42, 40, 41], and learning the dynamics of $E\times B$ plasmas [13, 14]. Nonlinear and high-dimensional data-driven models with control, often using deep reinforcement learning, have also been successfully applied in recent years to learn control policies for tokamak shape design [12] and instability control [49].
@@
In this work we focus on linear modeling and control of the injector circuits of the Helicity Injected Torus – Steady Inductive Upgrade (HIT-SIU), which are responsible for forming and sustaining a spheromak plasma inside the vacuum chamber. While these circuits exhibit a nonlinear coupling to the plasma, past work on the HIT-SI devices has shown that the nonlinear coil plasma system can be modeled as a linear system [33]. In addition, past work on tokamaks has shown that the nonlinear interaction between the plasma and the vertical control coils can be modeled and controlled as a linear system [31]. Further, we are interested in being able to easily modify an identified dynamics model without an expensive retraining phase, whether that be in terms of computational cost or accumulated training data. As such, the goal of this work is to implement the bagging optimized dynamic mode decomposition (BOP-DMD) [5, 47] to discover linear models for the interaction between spheromak plasmas formed in HIT-SIU, and the helicity injectors used to form and sustain the plasma. BOP-DMD has a number of advantages over standard DMD, including (i) being optimally robust to noise, (ii) providing uncertainty quantification metrics, and (iii) allowing for constraining the linear model to be stable [5]. BOP-DMD models are then paired with optimal control and estimation, in the form of linear quadratic Gaussian control (LQG), with the objective of controlling the current profiles of the flux and voltage coils on the helicity injectors.
@@
II HIT-SIU
The HIT-SIU experiment is studied using steady inductive helicity injection (SIHI) [27, 53] to form and sustain spheromak plasmas. Energy and helicity are injected into the plasma using a set of four injectors, each comprised of a set of two coils, a so called voltage coil and a flux coil, that link a semi-toroidal channel that connects to the main plasma volume. In each injector the flux coil injects, locally toroidal, magnetic flux, while the voltage coil acts as an air core transformer generating a voltage parallel to the flux-coil-generated field. The magnetic fields of the voltage and flux coils are orthogonal to one another such that there is no mutual inductance between the two coils when no plasma is present. However, all four injectors are part of a common manifold structure (Fig. 1), so the four voltage coils, and separately all four flux coils, are inductively coupled. Each coil is part of a larger resonant circuit comprised of an additional inductor and capacitor, which is driven by a switching power amplifier (SPA). The SPAs drive a three-state square wave with a high, low, and zero crossing time. An example power supply waveform is shown in Fig. 2. The injector circuit topology is shown on the left side of Fig. 3. While the voltage and flux circuits have the same topology, the values of each individual element differ between the two circuits. However, both circuits are tuned to the same resonant frequency.
@@ II HIT-SIU
II.1 Injector Circuit Control Methods:
Fast, and accurate control methods are critical for creating, and sustaining, high-performance plasmas in present-day and future plasma physics experiments. The shape of the voltage and current waveforms for the HIT-SIU voltage and flux coils can determine the kinds of instabilities induced in the spheromak, as well as the overall helicity injected into the spheromak. All coils of the HIT-SIU experiment were controlled using a GPU-based proportional integral derivative (PID) controller [38]. The PID algorithm parameterized the target waveforms to be driven as a sinusoid with an amplitude $V_{0}$ or $\Psi_{0}$, frequency $f_{\text{inj}}$, phase $\phi_{A},\phi_{B},\phi_{C},\phi_{D}$, and offset, and used these control parameters to drive the desired current waveform through the coil. If the frequency and amplitude of the waveform driven on each injector is the same, the expression for the rate of helicity injection is given by
@@ II.1 Injector Circuit Control Methods:
While the PID algorithm implemented in  [38] was a significant improvement upon the previous control scheme for the HIT-SI devices, further improvements can be made by shifting from model-free control to model-based control.
@@
Linear optimal control has been utilized across science and engineering to control complex systems [22]. Optimal linear control involves pairing an optimal state estimator, which reconstructs the full state of the system from limited noisy measurements, with a feedback controller, which calculates the optimal input to the system to track a desired trajectory. This optimal estimator, when constructed from a solution to the algebraic Riccati equation, is known as a Kalman filter, and the feedback controller, when also found as the solution to the Riccati equation, is called the linear quadratic regulator (LQR). These methods will pair together for LQG control. Both the Kalman filter and LQR rely on an underlying linear dynamics model of the system so that the current state and optimal gain to apply can be computed.
@@
Figure 3: Circuit topology for a voltage or a flux circuit for a single injector. This diagram includes the SPA (S1-S4), the series inductor (L1), a parallel capacitance (C), and a parallel inductance (L2) which represents the voltage or the flux coil. a) is an example of a flux or voltage circuit in vacuum, and b) is flux or voltage circuit when a plasma is present. Lp denotes the inductance of the plasma, Rp its resistance, and Mp the mutual inductance between the circuit and the plasma.
Figure 4: A representative equilibrium, with current gain of six, showing field lines linking (gray) and isolated (rainbow) from the injector volume.
III Analytic Model
@@
While there have been studies [30, 19, 20, 23] on the injector dynamics of the HIT-SI family of devices using MHD simulations and nonlinear reduced order models (ROMs), a linear model of the interaction between the plasma and the injector circuits for the purposes of control has not been developed. Before this study, the work of  [33], which derived a linear state space model of the injector circuit dynamics on HIT-SI3 and implemented a Kalman filter, was the only significant effort. To derive the linear model that will form the basis for the LQG control loop we seek to implement, we will begin with the circuit topology for the injector circuits as shown in Fig.  3. Treating the voltage across the capacitor, and the current through each coil as the states of our system, and then incorporating the mutual inductance between each flux or voltage coil, yields a 12-dimensional coupled system. This 12-dimensional system will be referred to as the vacuum circuit model. When a plasma is present one or more additional states need to be added to the model. In this work we choose to represent the plasma as a series LR circuit, which is coupled with a mutual inductance to each voltage or flux circuit, driven at or near its resonance. Therefore we elect to keep a fixed value for the inductance of the plasma. This yields a 13-dimensional coupled linear model. The full analytic vacuum model is available in Appendix A, and the analytic model for when a plasma is present is available in Appendix B.
There are of course many ways to represent the coupling between the injector circuits and the plasma as a linear system. For example, one may wish to add a capacitor to the circuit that represents the plasma. However, this would involve adding an additional voltage measurement to the state vector that represents the coupled injector-plasma system. By constructing a more complex circuit, a more accurate model of the coupled coil-plasma system may be derived, at the cost of additional model complexity and the inclusion of additional measurements. The methods proposed here are able to be adapted to any representation of the plasma as a linear circuit model.
@@
The values used in the state space model for the inductance of the various circuit elements were obtained by applying a Hilbert transform to the voltage and current waveforms from each element during an experimental vacuum shot. However, there is a slight difference between the capacitance used in the circuit model, and the listed capacitance for the flux circuits in HIT-SIU. The capacitance for each flux circuit in the model range from 94.08 to 123.96 microfarads. These values were chosen by inspection to achieve better matching between the phases of the flux coil current waveforms of the model, and the observed flux coil current waveforms during vacuum shots.
@@
This model does make some important assumptions. Specifically, we assume that when a plasma is present in the confinement volume, there is no inductive coupling between the plasma and the series inductor in each injector circuit. This assumption, which is well motivated as the series inductor is separated by a significant distance from the device, then allows us to conclude that the matrix that determines how actuation affects the dynamics of the injector circuits is constant between vacuum and plasma regimes. This can be observed from the structure of the $\mathbf{B}$ matrix shown in Appendix C. By using standard state space techniques for circuit analysis, it is clear that the power supply input will only directly change the state of the series inductor. Thus, the structure of the circuit is such that we can isolate differences between the vacuum and plasma regimes to the flux or voltage coils.
@@
Figure 6: The analytic linear model for the flux circuits when a plasma is present is simulated using a power supply waveform from the ninth shot on August 16th, 2022. The analytic model represents the spheromak as an inductor that is mutually coupled to the other flux circuits. To the left of the vertical blue line ($\approx 1$ ms), there is no plasma present, and to the right, a plasma is present.
III.1 Matlab and Simulink Model:
Developing control methods for hardware first requires a demonstration of performance in a digital environment. To that end, a model of the coupled flux circuits with full feedback control was developed in Matlab and Simulink. To verify the impedance of the various circuit elements, shot data was analyzed to capture the reactance and resistance of the elements, and to capture the strong and weak mutual inductance present between the injector circuits. Impedance is defined as
$\mathrm{Z}=R+iX.$ (2)
The real part of Equation 2 is the resistance of a circuit element, and the imaginary part of Equation 2 is the reactance. For an inductor, the inductance is related to the reactances by $\mathrm{Im(Z)}=i\omega L$. For a capacitor, the capacitance is related to the reactance by $\mathrm{Im(Z)}=\frac{1}{i\omega C}$  [2]. The resistance that is found for each circuit element is represented as a resistor in series with the element as shown in Fig.  3. Since the vacuum model can be understood as a special case of the plasma model with no mutual inductance between the plasma and the flux coils, we use the same impedance for the flux circuit elements across models. However, the value chosen for the impedance of the plasma in the plasma model, the resistance in the plasma, and the coupling between the plasma and the flux coils, were chosen by inspection, and intuition about the behavior of the plasma. To evaluate the performance of the analytic vacuum and plasma models with respect to capturing the dynamics present during an experimental shot, an SPA waveform from a shot is used as an input to the model, and the outputs of the model are evaluated against the flux circuit measurements from the experiment. The results of this test are shown in Fig. 5, and Fig. 6 for the vacuum and plasma models respectively. Fig. 7 shows that the amplitude and phase of the injector circuits are well matched by the analytic model. While the analytic model is able to capture much of the behavior of the flux circuits in vacuum, the plasma model struggles to capture the behavior of the flux circuits when a plasma is present.
@@ III.1 Matlab and Simulink Model:
With the derivation of a model that accurately captures the dynamics of interest in our system, the next step in the control scheme is to implement a Kalman filter. The Kalman filter uses a linear state space model of the injector circuit dynamics to estimate the state of the system from limited, and noisy measurements. We first demonstrate the performance of the Kalman filter when the underlying model is the analytic model for the injector circuits in vacuum, that was discussed previously in this section. Figure 8 demonstrates the performance of the Kalman filter when reconstructing the voltage across the capacitor in the first flux circuit. While the filter only has access to four noisy measurements of the flux coil currents, it is still able to accurately predict the value of the capacitor voltage throughout the shot.
@@
Finally, we contend with the controller: LQR minimizes a cost function $J$, that weighs the cost of control, with the value of producing an accurate reconstruction of the desired trajectory. These two costs are weighted by hyperparameters that can be tuned to prioritize accuracy of a specific state, and in this case was picked to be the current through the flux or voltage coil, as this trajectory is most closely linked to plasma performance. In this paradigm, a Kalman filter is used to predict from limited noisy measurements, the value of each state of the system at a given time. The LQR controller then makes a control decision to push the system to the desired state. The full control loop for controlling and estimating the injector circuits in vacuum is shown in Fig. 9.
@@
A user must decide what the desired trajectory for the system is. HIT-SIU has three primary operating regimes that differ in the relative temporal phasing of the current in each flux coil. For this simulation, we choose the case where all circuits have the same temporal phasing, which is the simplest case of operation on HIT-SIU. The desired current waveform is arrived at by simulating the analytic model of the circuits with a standard power supply waveform seen on experimental shots. This desired waveform is then subtracted off from the flux circuit predictions that are the outputs of the Kalman filter. The LQR controller gain is then applied to the full state estimation with the desired trajectory subtracted off, allowing for trajectory tracking, instead of fixed point tracking. A power supply waveform generated by the LQR controller is shown in Fig. 10. The action of the controller on the system produces a current profile through the flux circuit that closely matches the desired profile in Fig. 11.
@@
Figure 8: The Kalman filter has access to four noisy flux coil currents, and none of the other eight states, and is able to reconstruct the remaining states of the system based on the underlying model of the injector circuits in vacuum. Shown here is the reconstruction of the capacitor voltage from the first flux circuit.
@@
Figure 9: Full LQG feedback control loop in Simulink with the analytic model of the injector circuits in vacuum. The Kalman filter receives as input the states corresponding to the current through the flux coils, and from this measurement reconstructs the full state of the system. The desired flux coil current profiles are then subtracted off from the current state of flux coils, and fed into the LQR controller.
@@
Figure 10: Voltage waveform determined by the LQR controller to best track the desired flux coil waveform for the injector circuits in vacuum.
@@
Figure 11: Output of the flux coil very closely matches the desired current waveform for the injector circuits in vacuum.
IV Role of Data-Driven Models